A credit card company charges compound interest annually on any unpaid debts at 5%. Mr. Brian now has an outstanding debt of $4200. What will be his outstanding debt 3 years from now?
Answer #1

A=P\left(1+\frac{r}{n}\right)^{nt}

P=4200

r=5% = 0.05

n=1 for annual compound

t=3 years

A=4200\left(1+\frac{0.05}{1}\right)^{1\cdot 3}

A=4200\left(1+0.05\right)^3

A=4200\left(1.05\right)^3

A=4200\cdot \:1.157625

{\color{Red} A=4862.025}...............outstanding debt 3 years from now